OpenStudy (anonymous):

If A = {positive even integers} and B = {positive odd integers}, what is A union B? {positive integers} {integers} {0} { } Question 31 (Multiple Choice Worth 2 points) [2.02] Solve for x: -3(x - 3) = 21 x = -10 x = 10 x = -4 x = 4

hero (hero):

@mcsmarty, for the first question, which choice do you believe is correct?

OpenStudy (anonymous):

integers so b

hero (hero):

Integers would include both positive and negative numbers. However sets A and B both include positive numbers that are either even or odd.

OpenStudy (anonymous):

so its positive

OpenStudy (anonymous):

a

hero (hero):

Positive Integers would be the correct choice because positive integers include all integers that are either even or odd and greater than zero.

OpenStudy (anonymous):

thanks nect one

hero (hero):

Have you attempted solving for x yet?

OpenStudy (anonymous):

yea I thinks its -10

hero (hero):

Did you check your answer? Do you know how to "check" after solving?

OpenStudy (anonymous):

ya u plug it in but they all came up with weird numbers that's were I need help

hero (hero):

Basically, there are two ways to solve the problem. You either use algebra or you use the check feature.

hero (hero):

You're already given the answer choices. If you select a choice and it doesn't pass the "check" move on to the next number.

OpenStudy (anonymous):

its either -10 or or -4 right?

hero (hero):

Well yes it is one or the other, but using the "check" feature, you should be able to figure out which of the two is correct.

OpenStudy (anonymous):

check feature on what open study or paper or internet

hero (hero):

For example, to check if it is -10, perform the following process: Replace x with -10, then simplify: -3(-10 - 3) = 21 -3(-13) = 21 39 = 21 FALSE

hero (hero):

So based on that checking process, we conclude that -10 is not the correct choice since both sides are not equal.

hero (hero):

The correct x leads to both sides being equal

OpenStudy (anonymous):

-4

hero (hero):

Yes -4 is the correct choice, but the question is, do you know how to check and do you know how to solve the equation algebraically?

OpenStudy (anonymous):

yea I was doing something wrong I got it now